# On $[1,2]$-Domination in Interval and Circle Graphs

### v2

@@ 2 Algorithm for interval graphs
Now, we are ready to present our algorithm (Algorithm 1) for computing $\gamma_{[1,2]}(G)$, where $G$ is an interval graph. Note that if $12\in E$, then the case (iv.b) of Lemma 7 holds and so $\gamma^{1}_{[1,2]}(1,2:2)<\gamma^{1}_{[1,2]}(1,2:2,1)$. By considering Lemmas 2, 3, 5, 6, 7 and 8 we see that if we use $\gamma^{1}_{[1,2]}(1,2:2,1)$ for computing a value, then we also use $\gamma^{1}_{[1,2]}(1,2:2)$ for computing that value. Hence, in Line 12 we set $\gamma^{1}_{[1,2]}(1,2:2,1)$ to $\infty$. We also use this note in Line 20 for computing $\gamma^{1}_{[1,2]}(k,i:i,j)$.
@@
5:$\gamma^{0}_{[1,2]}(1)\leftarrow\infty$
6:If $12\in E$, then $\gamma^{0}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(2)\leftarrow\infty$
7:$\gamma^{0}_{[1,2]}(1,2)\leftarrow\infty$
8:If $12\in E$, then $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow\infty$
9:$\gamma^{1}_{[1,2]}(1)\leftarrow 1$
10:If $12\in E$, then $\gamma^{1}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{1}_{[1,2]}(2)\leftarrow 2$
11:If $12\in E$, then $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow 1$, otherwise, $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow\infty$
$\gamma^{1}_{[1,2]}(1,2:2,1)\leftarrow\infty$
$\gamma^{11}_{[1,2]}(1,3:3,2,1)\leftarrow 3$
@@
17:         Compute $\gamma^{1}_{[1,2]}(j,i:i)$ using Lemma 6.
@@
20:         Compute $\gamma^{0}_{[1,2]}(k,i:j)$ using Lemma 3.
21:         Compute $\gamma^{1}_{[1,2]}(k,i:i,j)$ using Lemma 7.
@@
24:         Compute $\gamma^{1}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 8.
25:         Compute $\gamma^{11}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 9.
@@
28:     Compute $\gamma^{1}_{[1,2]}(i)$ using Lemma 5.
@@
30:return $\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$
Theorem 1.
@@
Proof.
Let $(1,2,\ldots,n)$ be a numbering of vertices of $G$ computed in Line 1. This numbering of vertices of $G=(V,E)$ can be computed in $O(|V|+|E|)$ time [12]. We need $O(n^{3})$ time to compute $\mathtt{low}(i)$ for all $i\in[1,n]$, $\mathtt{low}(j,i)$ for all $1\leq j<i\leq n$ and $\mathtt{maxlow}(i)$ for all $i\in[1,n]$ in Lines 2-4. It is easy to compute the values $\gamma^{0}_{[1,2]}(1)$, $\gamma^{0}_{[1,2]}(2)$, $\gamma^{0}_{[1,2]}(1,2)$, $\gamma^{0}_{[1,2]}(1,2:1)$, $\gamma^{1}_{[1,2]}(1)$, $\gamma^{1}_{[1,2]}(2)$, $\gamma^{1}_{[1,2]}(1,2:2)$, $\gamma^{1}_{[1,2]}(1,2:2,1)$ and $\gamma^{11}_{[1,2]}(1,3:3,2,1)$ as seen in Lines 5-13 in $O(1)$ time. Let $i\in[3,n]$, $j\in[1,i)$, $k\in[1,j]$, $k^{\prime}\in[2,j)$ and $l\in[1,k^{\prime})$. By Lemmas 2, 3, 5, 6, 7, 8 and 9, compute $\gamma^{0}_{[1,2]}(i)$, $\gamma^{0}_{[1,2]}(j,i)$, $\gamma^{0}_{[1,2]}(k,i:j)$, $\gamma^{1}_{[1,2]}(i)$, $\gamma^{1}_{[1,2]}(j,i:i)$, $\gamma^{1}_{[1,2]}(k,i:i,j)$, $\gamma^{1}_{[1,2]}(k,i:i,j,k^{\prime})$ and $\gamma^{11}_{[1,2]}(k,i:i,j,k^{\prime})$ in Lines 15-29 We see that the running time of Lines 14-29 is $O(n^{4})$. Hence, the running time of Algorithm 1 is $O(n^{4})$. Clearly, $\gamma_{[1,2]}(G)=\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$. This completes the proof of the theorem. ∎
@@ 3 $[1,2]$-domination in Circle Graphs
In this section, we prove hardness results for a $[1,2]$-domination problem in circle graphs. We start with the main result of this section.
Theorem 2.
$[1,2]$-domination problem is $NP$-complete in circle graphs.
Proof.
Clearly $[1,2]$-domination problem is in $NP$. To prove that the problem is $NP$-hard, We shall reduce (in polynomial time) the $3SAT$ problem to the problem of finding a $[1,2]$-dominating set. Let $X=\{x_{1},x_{2},\dots,x_{n}\}$ be the set of boolean variable and $C=\{C_{1},C_{2},\dots,C_{m}\}$ be the set of clauses in an arbitrary instance of $3SAT$, like $I$. In the first part, we construct a circle graph $G(I)$ based on the instance $I$ and in the second part, we choose an integer $k$ and show $G(I)$ has $[1,2]$-dominating set of size $k$ if and only if $I$ is satisfiable. Let us describe the construction of $G(I)$ produced in reduction from $3SAT$ to $[1,2]$-domination problem. Without loss of generality, we assume that every variable appears once in each clause. We choose an arbitrary point of the circle as the origin. The circle graph $G(I)$ is defined as follows:
@@
    We divide the circle into $m$ disjoint open intervals $[s_{j},s^{\prime}_{j}]$ for $1\leq j\leq m$, corresponding each clause,
    Figure 10: Intervals corresponding each clause
@@
    We create two chords $t_{j}^{l}$ and $f_{j}^{l}$ associated with each literal $l$, $1\leq l\leq 3$, that appears in clause $C_{j}$. Including the chord $t_{j}^{l}$ in $[1,2]$-dominating set will correspond to setting the associated literal to true and including the $f_{j}^{l}$ chord in $[1,2]$-dominating set will correspond to setting the associated literal to false,
@@
    For each literal in clause $C_{j}$ we create a pair of chords $a_{j}^{l}$,
@@
    For each literal in clause $C_{j}$ we create a pair of chords $p_{j}^{l}$. The purpose of each such pair is to dominate them by including the chords $t_{j}^{l}$ or $f_{j}^{l}$,
@@
    For each literal in clause $C_{j}$, the interval between the endpoint of $t_{j}^{l}$ and the endpoint of the first chords of $a_{j}^{l}$ is called the truth interval, and the interval between the right point of the second chord of $p_{j}^{l}$ and the right point of $f_{j}^{l}$ is called false interval for literal $l$ of clause $j$. (See Figure 11)
@@
    For each clause $C_{j}$, we create three pairs of chords of types $u$ and $w$. Based on the endpoints of $a_{j}^{l}$ pairs in clause $C_{j}$, we have pairs $(u_{12},w_{12})$,$(u_{13},w_{13})$ and, $(u_{23},w_{23})$. Also, we have two guard chords, $g_{1}$ and $g_{2}$, connected to two the claw graphs centered at $c_{1}$ and $c_{2}$, respectively. The purpose of each will be discussed later.
    • The chord $u_{13}$ incidence with only the pair $a_{j}^{1}$ and $w_{13}$ incidence with only the pair $a_{j}^{3}$,
    • The left point of $u_{12}$ is before pair $a_{j}^{1}$ and endpoint of it is before the pair $a_{j}^{3}$. So, the chord $u_{12}$ incidence with the pairs $a_{j}^{1}$ and $a_{j}^{2}$. The chord $w_{12}$ start before the endpoint $u_{13}$ and after crossing the chord $u_{13}$, the pairs $a_{j}^{2}$, and the chord $w_{13}$ will be finished,
    • The left point of $u_{23}$ is after pair $a_{j}^{1}$ and endpoint of it is after the pair $a_{j}^{3}$. So, the chord $u_{23}$ incidence with the pairs $a_{j}^{2}$ and $a_{j}^{3}$. The chord $w_{23}$ start before the endpoint $u_{13}$ and also before the start point of $w_{12}$. After crossing the chord $u_{13}$, the pairs $a_{j}^{2}$, and the chords $w_{12},w_{13}$ will be finished,
    • The guard chord $g_{1}$, from the left endpoint is connected to a chord $c_{1}$ which three independent chords cross it. The purpose of these independent chords is to force that $c_{1}$ contains in any minimum $[1,2]$-dominating set. So, the guard chord $g_{1}$ is always dominated by $c_{1}$ once. The right endpoint of $g_{1}$ will be finished after the start point of $u_{23}$. So, $g_{1}$ will incidence with the chords $u_{12},u_{13}$, and $u_{23}$,
    • The left endpoint of $g_{2}$ will start before the left endpoint of $u_{12}$ and finish after the right endpoint of $w_{13}$. So, $g_{2}$ will incidence with the chords $u_{12},u_{23}$, and $w_{13}$. The guard chord $g_{2}$, from the right endpoint, is connected to a chord $c_{2}$ which three independent chords cross it. Similar to chord $g_{1}$, the purpose of these independent chords is to force that $c_{2}$ contains in any minimum $[1,2]$-dominating set. So, the guard chord $g_{2}$ is always dominated by $c_{2}$ once.
@@
The key idea of our proof is as follows. Suppose that $C_{j}$ is satisfied by some assignment of truth values to a variable. Then we can construct a $[1,2]$-dominating set that contains at least one chord $t_{j}^{l}$ corresponding to a true literal in $C_{j}$ and a pair of $u$ and $w$ chords associated with $C_{j}$. The chord $t_{j}^{l}$ dominates the pairs $a_{j}^{l}$ while the other pairs $a_{j}^{l}$ and $(u,w)$’s are dominated by an appropriate pair of $(u,w)$. To ensure that exactly one of the chords $t_{j}^{l}$ and $f_{j}^{l}$ associated with literal $l$ in clause $j$ appears in any $[1,2]$-dominating set, a pair of independent chords $p_{j}^{l}$ associated with the literal $l$ adjacent to both $t_{j}^{l}$ and $f_{j}^{l}$ and no other chords. In other words, since the independent chords $p_{j}^{l}$’s start from the left endpoint of $f_{j}^{l}$ and the right endpoint $t_{j}^{l}$, to dominate them, $t_{j}^{l}$ or $f_{j}^{l}$ or both of chords in $p_{j}^{l}$ should be select. Later we will argue that any minimum $[1,2]$-dominating set must contain exactly one of chords $t_{j}^{l}$ or $f_{j}^{l}$ otherwise both of $p_{j}^{l}$ must be selected that leads to a larger dominating set.
The clause $C_{j}$ is satisfied if and only if at least one of three literals involved in $C_{j}$ is true. As we mentioned, this will correspond to at least one of the chords $t_{j}^{l}$ appearing in the dominating set. So this chord will dominate $t_{j}^{l}$, $f_{j}^{l}$, and both pairs $a_{j}^{l}$ and $p_{j}^{l}$. To dominate the remaining pairs $a_{j}$ associated with other literals in $C_{j}$ we use six chords of type $u$ and $w$ as follows. $u_{13}$ and $w_{13}$ cross the pairs $a_{j}^{1}$ and $a_{j}^{3}$, respectively. If the second literal is true that leads $t_{j}^{2}$ be in $[1,2]$-dominating set, two pairs $a_{j}^{1}$ and $a_{j}^{3}$ remain undominated, so selecting $u_{13}$ and $w_{13}$ dominate them.
@@
In the above construction, we obtain $m$ pairwise disjoint sections, each associated with one clause. In the second part of our construction, we use connection chords that force a variable to have the same truth value throughout all clauses. It is known as the consistency condition. we create connection chords $tt$, $ff$, $tf$, and $ft$ that connect the clauses as follows. So, the consistency of truth values of literals throughout all clauses will be preserved.
@@
Figure 13 shows all chords associated with a two clause and four variables.
@@
Let $k=7m+2t$ where $m$ is the number of clauses and $t$ is the total number of repetitions of variables in clauses. Theorem 2 directly follows from the following result.
Lemma 10.
@@
Proof.
Given a satisfying truth assignment for $I$, we show how to construct a $[1,2]$-dominating set $D$ of size $k=7m+2t$. Let us consider that the clause $C_{j}$ contains three literals corresponding to variables $x_{i},x_{k}$, and $x_{l}$ where $i\leq k\leq l$. We assume the variable $x_{i}$ appears in the first literal and $x_{k},x_{l}$ are the second and third literals in $C_{j}$, respectively. Since $I$ is satisfied, at least one of the literals $x_{i},x_{k}$, and $x_{l}$ must be true. Let’s assume without loss of generality that the only first literal of $C_{j}$ is true and the second and third literals are false according to the assignment. So we have included $t_{j}^{l},f_{j}^{2}$, and $f_{j}^{3}$ in $D$. This ensures that all the pairs $p_{1},p_{2}$, and $p_{3}$ have been dominated. Also, the pair $a_{j}^{1}$ is dominated, but the pairs $a_{j}^{2}$ and $a_{j}^{3}$ are not dominated. To dominate them and all pairs $(u,w)$ in this clause, we put $u_{23},w_{23}$ to $D$. So to dominate this clause, we need at least five chords. Also the chords $c_{1}$ and $c_{2}$ must be put in $D$ to dominate the three independent chords crossing them.
Generally, for each literal $l$ in clause $C_{j}$ that is true in the assignment of $I$, we include in $D$ the chord(s) $t_{j}^{l}$ where $l$ is the literal number in clause $C_{j}$. Otherwise, we include in $D$ the chord(s) $f_{j}^{l}$. The chords $t_{j}^{l}$ dominate the pair $a_{j}^{l}$. So, to dominate the remaining $a_{j}$ pairs and the pairs $(u,w)$ in $C_{j}$, we should choose at least one of $(u,w)$ pairs. Also the chords $c_{1}$ and $c_{2}$ must be put in $D$.
For each chord $tt,tf,ft$, and $ff$, the chord $s$ connected to them put in dominating set so all these chords are dominated once. otherwise, we must choose three independent chords connected to $s$ the leads $s$ be dominated three times and contradiction. So for each chord $tt,tf,ft$, and $ff$, we need one chord in dominating set. According to the construction for each variable repeated in two clauses, we have two connection chords. If the total number of variable repartition is $t$, we include $2t$ chords in $D$. It is clear that $D$ is a $[1,2]$-dominating set.
Lemma 11.

### v4

@@ 2 Algorithm for interval graphs
Now, we are ready to present our algorithm (Algorithm 1) for computing $\gamma_{[1,2]}(G)$, where $G$ is an interval graph. Note that if $12\in E$, then the case (iv.b) of Lemma 2.10 holds and so $\gamma^{1}_{[1,2]}(1,2:2)<\gamma^{1}_{[1,2]}(1,2:2,1)$. By considering Lemmas 4, 2.2, 2.6, 2.8, 2.10 and 2.12 we see that if we use $\gamma^{1}_{[1,2]}(1,2:2,1)$ for computing a value, then we also use $\gamma^{1}_{[1,2]}(1,2:2)$ for computing that value. Hence, in Line $12$ we set $\gamma^{1}_{[1,2]}(1,2:2,1)$ to $\infty$. We also use this note in Line $21$ for computing $\gamma^{1}_{[1,2]}(k,i:i,j)$.
@@
5:$\gamma^{0}_{[1,2]}(1)\leftarrow\infty$.
6:If $12\in E$, then $\gamma^{0}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(2)\leftarrow\infty$.
7:$\gamma^{0}_{[1,2]}(1,2)\leftarrow\infty$.
8:If $12\in E$, then $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow\infty$.
9:$\gamma^{1}_{[1,2]}(1)\leftarrow 1$.
10:If $12\in E$, then $\gamma^{1}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{1}_{[1,2]}(2)\leftarrow 2$.
11:If $12\in E$, then $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow 1$, otherwise, $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow\infty$.
$\gamma^{1}_{[1,2]}(1,2:2,1)\leftarrow\infty$.
$\gamma^{11}_{[1,2]}(1,3:3,2,1)\leftarrow 3$.
@@
17:         Compute $\gamma^{1}_{[1,2]}(j,i:i)$ using Lemma 2.8.
@@
20:         Compute $\gamma^{0}_{[1,2]}(k,i:j)$ using Lemma 2.2.
21:         Compute $\gamma^{1}_{[1,2]}(k,i:i,j)$ using Lemma 2.10.
@@
24:         Compute $\gamma^{1}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 2.12.
25:         Compute $\gamma^{11}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 2.13.
@@
28:     Compute $\gamma^{1}_{[1,2]}(i)$ using Lemma 2.6.
@@
30:return $\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$.
Theorem 2.14.
@@
Proof 2.15.
Let $(1,2,\ldots,n)$ be a numbering of vertices of $G$ computed in Line 1. This numbering of vertices of $G=(V,E)$ can be computed in $O(|V|+|E|)$ time Ramalingam and Rangan (1988). We need $O(n^{3})$ time to compute $\mathtt{low}(i)$ for all $i\in[1,n]$, $\mathtt{low}(j,i)$ for all $1\leq j<i\leq n$ and $\mathtt{maxlow}(i)$ for all $i\in[1,n]$ in Lines 2-4. It is easy to compute the values $\gamma^{0}_{[1,2]}(1)$, $\gamma^{0}_{[1,2]}(2)$, $\gamma^{0}_{[1,2]}(1,2)$, $\gamma^{0}_{[1,2]}(1,2:1)$, $\gamma^{1}_{[1,2]}(1)$, $\gamma^{1}_{[1,2]}(2)$, $\gamma^{1}_{[1,2]}(1,2:2)$, $\gamma^{1}_{[1,2]}(1,2:2,1)$ and $\gamma^{11}_{[1,2]}(1,3:3,2,1)$ as seen in Lines 5-13 in $O(1)$ time. Let $i\in[3,n]$, $j\in[1,i)$, $k\in[1,j]$, $k^{\prime}\in[2,j)$ and $l\in[1,k^{\prime})$. By Lemmas 4, 2.2, 2.6, 2.8, 2.10, 2.12 and 2.13, compute $\gamma^{0}_{[1,2]}(i)$, $\gamma^{0}_{[1,2]}(j,i)$, $\gamma^{0}_{[1,2]}(k,i:j)$, $\gamma^{1}_{[1,2]}(i)$, $\gamma^{1}_{[1,2]}(j,i:i)$, $\gamma^{1}_{[1,2]}(k,i:i,j)$, $\gamma^{1}_{[1,2]}(k,i:i,j,k^{\prime})$ and $\gamma^{11}_{[1,2]}(k,i:i,j,k^{\prime})$ in Lines 15-29 We see that the running time of Lines 14-29 is $O(n^{4})$. Hence, the running time of Algorithm 1 is $O(n^{4})$. Clearly, $\gamma_{[1,2]}(G)=\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$. This completes the proof of the theorem.
@@ 3 $[1,2]$-domination in Circle Graphs
In this section, we prove hardness results for the $[1,2]$-domination problem in circle graphs. We start with the main result of this section.
Theorem 3.16.
The $[1,2]$-domination problem is $NP$-complete in circle graphs.
Proof 3.17.
Clearly, the $[1,2]$-domination problem is in $NP$. To prove that the problem is $NP$-hard, We shall reduce (in polynomial time) the $3SAT$ problem to the problem of finding a $[1,2]$-dominating set. Let $X=\{x_{1},x_{2},\dots,x_{n}\}$ be the set of boolean variable and $C=\{C_{1},C_{2},\dots,C_{m}\}$ be the set of clauses in an arbitrary instance of $3SAT$, like $I$. In the first part, we construct a circle graph $G(I)$ based on the instance $I$, and in the second part, we choose an integer $k$ and show $G(I)$ has a $[1,2]$-dominating set of size $k$ if and only if $I$ is satisfiable. Let us describe the construction of $G(I)$ produced in the reduction from $3SAT$ to the $[1,2]$-domination problem. Without loss of generality, we assume that every variable appears once in each clause. We choose an arbitrary point of the circle as the origin. The circle graph $G(I)$ is defined as follows:
@@
    We divide the circle into $m$ disjoint open intervals $[s_{j},s^{\prime}_{j}]$ for $1\leq j\leq m$, corresponding to each clause,
    Figure 10: Intervals corresponding to each clause
@@
    We create two chords $t_{j}^{l}$ and $f_{j}^{l}$ associated with each literal $l$, $1\leq l\leq 3$, that appears in clause $C_{j}$. Including the chord $t_{j}^{l}$ in a $[1,2]$-dominating set will correspond to setting the associated literal to true and including the $f_{j}^{l}$ chord in a $[1,2]$-dominating set will correspond to setting the associated literal to false,
@@
    For each literal in clause $C_{j}$, we create a pair of chords $a_{j}^{l}$,
@@
    For each literal in clause $C_{j}$, we create a pair of chords $p_{j}^{l}$. The purpose of each such pair is to dominate them by including the chords $t_{j}^{l}$ or $f_{j}^{l}$,
@@
    For each literal in clause $C_{j}$, the interval between the endpoint of $t_{j}^{l}$ and the endpoint of the first chord of $a_{j}^{l}$ is called the truth interval, and the interval between the right endpoint of the second chord of $p_{j}^{l}$ and the right endpoint of $f_{j}^{l}$ is called false interval for literal $l$ of clause $C_{j}$. (See Figure 11)
@@
    For each clause $C_{j}$, we create three pairs of chords of types $u$ and $w$. Based on the endpoints of $a_{j}^{l}$ pairs in clause $C_{j}$, we have pairs $(u_{12},w_{12})$,$(u_{13},w_{13})$ and $(u_{23},w_{23})$. Also, we have two guard chords, $g_{1}$ and $g_{2}$, connected to the two claw graphs centered at $c_{1}$ and $c_{2}$, respectively. The purpose of each will be discussed later.
    • The chord $u_{13}$ is incident with only the pair $a_{j}^{1}$ and $w_{13}$ is incident with only the pair $a_{j}^{3}$,
    • The left point of $u_{12}$ is before pair $a_{j}^{1}$ and endpoint of it is before the pair $a_{j}^{3}$. So, the chord $u_{12}$ is incident with the pairs $a_{j}^{1}$ and $a_{j}^{2}$. The chord $w_{12}$ starts before the endpoint $u_{13}$ and after crossing the chord $u_{13}$, the pairs $a_{j}^{2}$, and the chord $w_{13}$ will be finished,
    • The left point of $u_{23}$ is after pair $a_{j}^{1}$ and endpoint of it is after the pair $a_{j}^{3}$. So, the chord $u_{23}$ is incident with the pairs $a_{j}^{2}$ and $a_{j}^{3}$. The chord $w_{23}$ starts before the endpoint $u_{13}$ and also before the start point of $w_{12}$. After crossing the chord $u_{13}$, the pairs $a_{j}^{2}$, and the chords $w_{12},w_{13}$ will be finished,
    • The guard chord $g_{1}$, from the left endpoint is connected to a chord $c_{1}$ which three independent chords cross. The purpose of these independent chords is to ensure that $c_{1}$ is included in any minimum $[1,2]$-dominating set. Therefore, the guard chord $g_{1}$ is always dominated by $c_{1}$ once. The right endpoint of $g_{1}$ will be finished after the start point of $u_{23}$. So, $g_{1}$ be incident with the chords $u_{12},u_{13}$, and $u_{23}$,
    • The left endpoint of $g_{2}$ starts before the left endpoint of $u_{12}$ and finishes after the right endpoint of $w_{13}$. Thus, $g_{2}$ be incidence with the chords $u_{12},u_{23}$, and $w_{13}$. The guard chord $g_{2}$, from the right endpoint, is connected to a chord $c_{2}$ which is crossed by three independent chords. Similar to chord $g_{1}$, the purpose of these independent chords is to ensure that $c_{2}$ is included in any minimum $[1,2]$-dominating set. So, the guard chord $g_{2}$ is always dominated by $c_{2}$ once.
@@
The key idea of our proof is as follows. Suppose that $C_{j}$ is satisfied by some assignment of truth values to the variables. Then we can construct a $[1,2]$-dominating set that contains at least one chord $t_{j}^{l}$ corresponding to a true literal in $C_{j}$ and a pair of $u$ and $w$ chords associated with $C_{j}$. The chord $t_{j}^{l}$ dominates the pairs $a_{j}^{l}$ while the other pairs $a_{j}^{l}$ and $(u,w)$’s are dominated by an appropriate pair of $(u,w)$. To ensure that exactly one of the chords $t_{j}^{l}$ and $f_{j}^{l}$ associated with literal $l$ in clause $j$ appears in any $[1,2]$-dominating set, a pair of independent chords $p_{j}^{l}$ associated with literal $l$ is adjacent to both $t_{j}^{l}$ and $f_{j}^{l}$ and no other chords. In other words, since the independent chords $p_{j}^{l}$’s start from the left endpoint of $f_{j}^{l}$ and end at the right endpoint $t_{j}^{l}$, to dominate them, $t_{j}^{l}$ or $f_{j}^{l}$ or both chords in $p_{j}^{l}$ must be select. Later we will argue that any minimum $[1,2]$-dominating set must contain exactly one of the chords $t_{j}^{l}$ or $f_{j}^{l}$ otherwise both of $p_{j}^{l}$ must be selected, which would lead to a larger dominating set.
The clause $C_{j}$ is satisfied if and only if at least one of three literals involved in $C_{j}$ is true. As we mentioned, this will correspond to at least one of the chords $t_{j}^{l}$ appearing in the dominating set. So this chord will dominate $t_{j}^{l}$, $f_{j}^{l}$, and both pairs $a_{j}^{l}$ and $p_{j}^{l}$. To dominate the remaining pairs $a_{j}$ associated with other literals in $C_{j}$ we use six chords of type $u$ and $w$ as follows. $u_{13}$ and $w_{13}$ cross the pairs $a_{j}^{1}$ and $a_{j}^{3}$, respectively. If the second literal is true that leads to $t_{j}^{2}$ being in the $[1,2]$-dominating set, two pairs $a_{j}^{1}$ and $a_{j}^{3}$ remain undominated. Hence, selecting $u_{13}$ and $w_{13}$ dominates them.
@@
In the above construction, we obtain $m$ pairwise disjoint sections, each associated with one clause. In the second part of our construction, we use connection chords that force a variable to have the same truth value throughout all clauses. This is known as the consistency condition. we create connection chords $tt$, $ff$, $tf$, and $ft$ that connect the clauses as follows. Thus, the consistency of the truth values of literals throughout all clauses will be preserved.
@@
Figure 13 shows all chords associated with two clauses and four variables.
@@
Let $k=7m+2t$ where $m$ is the number of clauses and $t$ is the total number of repetitions of variables in clauses. Theorem 3.16 directly follows from the following result.
Lemma 3.18.
@@
Proof 3.19.
Given a satisfying truth assignment for $I$, we show how to construct a $[1,2]$-dominating set $D$ of size $k=7m+2t$. Let us consider that the clause $C_{j}$ contains three literals corresponding to variables $x_{i},x_{k}$, and $x_{l}$ where $i\leq k\leq l$. We assume the variable $x_{i}$ appears in the first literal and $x_{k},x_{l}$ are the second and third literals in $C_{j}$, respectively. Since $I$ is satisfied, at least one of the literals $x_{i},x_{k}$, and $x_{l}$ must be true. Assume without loss of generality, that the first literal of $C_{j}$ is true and the second and third literals are false according to the assignment. So we have included $t_{j}^{l},f_{j}^{2}$, and $f_{j}^{3}$ in $D$. This ensures that all the pairs $p_{1},p_{2}$, and $p_{3}$ have been dominated. Also, the pair $a_{j}^{1}$ is dominated, but the pairs $a_{j}^{2}$ and $a_{j}^{3}$ are not dominated. To dominate them and all pairs $(u,w)$ in this clause, we put $u_{23},w_{23}$ to $D$. Thus, to dominate this clause, we need at least five chords. Additionally, the chords $c_{1}$ and $c_{2}$ must be put in $D$ to dominate the three independent chords crossing them.
Generally, for each literal $l$ in clause $C_{j}$ that is true in the assignment of $I$, we include the chord(s) $t_{j}^{l}$ in $D$ where $l$ is the literal number in clause $C_{j}$. Otherwise, we include in $D$ the chord(s) $f_{j}^{l}$. The chords $t_{j}^{l}$ dominate the pair $a_{j}^{l}$. To dominate the remaining $a_{j}$ pairs and the pairs $(u,w)$ in $C_{j}$, at least one of $(u,w)$ pairs must be chosen. Additionally, the chords $c_{1}$ and $c_{2}$ must be included in $D$.
For each chord $tt,tf,ft$, and $ff$, the chord $s$ connected to them is put in dominating set so all these chords are dominated once. otherwise, we must choose three independent chords connected to $s$, which would lead to $s$ being dominated three times, resulting in a contradiction. Therefore, for each chord $tt,tf,ft$, and $ff$, we need one chord in the dominating set. According to the construction, for each variable repeated in two clauses, we have two connection chords. If the total number of variable repartition is $t$, we include $2t$ chords in $D$. It is clear that $D$ is a $[1,2]$-dominating set.
Lemma 3.20.